# Statistical Learning for Heterogeneous Treatment Effects: Pretraining, Prognosis, and Prediction

### v1

@@ 5 Discussion
A number of important extensions of our approach are left option. The proposed forms of supervised pretraining used in this work is demonstrated for settings where researchers have access to a single dataset. However, in many settings it might be valuable to pretrain CATE models on an additional external dataset. Our approach can be simply extented to those setting as it is possible to pretrain the models for $m(X_{i})$, $e(X_{i})$ and $\tau(X_{i})$ using an external dataset. Craig et al. (2024) demonstrate that training a predictive model across multiple tumor types and then transferring knowledge from this pan-cancer model to a subtype-specific model (e.g., for breast cancer) can lead to improved predictive accuracy. Similarly, in cancer trials that are further stratified by biomarkers, it might be beneficial to pretrain a model on the total trial population (intention-to-treat) and transfer this information into a CATE model fit on a subgroup that is enriched for a given biomarker. This can help overcome some settings with limited sample size, high dimensions, and low signal.
In follow-up work, we aim to demonstrate the utility of this approach in real-world data from clinical settings, with a focus on applications in oncology.
@@ Acknowledgments
We are grateful to Erin Craig, Trevor Hastie, Guido Imbens, Ying Lu, Jonathan Taylor, Lu Tian and Stefan Wager as well as seminar participants at Stanford University for valuable comments and feedback.

### v2

@@ 5 Discussion
A number of extensions of our approach are left open. The proposed forms of supervised pretraining used in this work are demonstrated for settings where researchers have access to a single dataset. However, in many settings, it might be valuable to pretrain CATE models on an additional external dataset. Our approach can be extended to those settings as it is possible to pretrain the models for $m(X_{i})$, $e(X_{i})$, and $\tau(X_{i})$ using an external dataset. Similarly, researchers may have access to a large dataset spanning multiple cancer types that share some common features. In this case, the $R$-Learner can first be trained on the full dataset, with penalties from each prediction task transferred to refine a second, cancer-specific $R$-Learner (e.g., for breast cancer or subtypes defined by biomarkers such as HER2). Pretraining can help obtain targeted insights, a task often made difficult by low signal, small sample sizes, and high dimensionality. Finally, our approach to incorporating side information relied on a covariate support set. Exploring alternative strategies that directly leverage coefficient estimates or other model components, such as baseline risk probabilities (perhaps via Bayesian approaches (Hill, 2011)), presents interesting directions for future work.
@@ Acknowledgments
We are grateful to Erin Craig, Trevor Hastie, Guido Imbens, Ying Lu, Jonathan Taylor, Lu Tian, and Stefan Wager, as well as seminar participants at Stanford University, for valuable comments and feedback.
@@
Declaration of interest statement
The authors declare no conflict of interest.
@@ Appendix A Supplement
Figure 10: Pretraining of the DR-Learner. The barplots shows model the performance in terms of mean squared error (MSE) when the CATE function in DR-Learner is pretrained using the adaptive lasso and side information from the controls through $\mu_{0}(x)$, the treated units through $\mu_{1}(x)$ or both. The MSE is normalized by the $DR$-lasso with no pretraining ($\alpha$ = 1) for each support size. The experiment is conducted using the DGP described subsection 3.2, with $p=60$, $n=500$ and $SNR=2$.